a die is rolled ten times.find the number of ways so that outcomes always contain 1,2,3.


please explain solution as well..

Grade:12

2 Answers

Sudheesh Singanamalla
13 years ago

Dear Kunal,

the number of times the number 1,2,3 appear when the dice is rolled once is 1/3

so number of times it is possible to get 1,2,3 when dice is rolled 10 times is 1/3^ 10

1/59049 = 0.000016935

it is raised to the power 10 because when 10 dice rolled, the number or possibilities = 6^10 = 60466176
and the number of possibilities of getting 1,2,3 are 103 approximately.

approve if you are pleased with the answer

Vishwajeet Singh
13 years ago

total no. of possibilities=6^10

You have got 10 places,

choose any 3 places by 10 C 3 ways=120 and fix 1 2 & 3 there...

7 places are left, they can have any of numbers, your conditions will be satisfied.

so the ANSWER is 120*(6^7)